what are the x and y coordinates for the equation y=3x-4
thanks a lot for helping me!

Answer by Abbey(339)

x and y coordinates are determined by the values you plug into the equation. The y is the dependant variable and the x is the independant variable. In other words, the value for y is based on what you select for x.
If you select x=0, then in this equation, y will be -4, you would say the x and y coordinates are (0,-4)
y=3x-4
y=-3(0)-4
y=0-4
y=-4
It is completely up to you which number you want to plug in for the x value.
If you chose x=1, then y would be -1 and your coordinate would be (1,-1)
because
y=3x-4
y=3(1)-4
y=3-4
y=-1